OpenStudy (anonymous):

A professor asked the students in his class to rate the guest lecturer. The lecturer received a rating of Excellent from 7 of the students, while 14 students rated the lecturer as Average, and 4 believed the lecturer was Poor. What percent of the students thought the lecturer was Excellent?

OpenStudy (valpey):

7/(7+14+4)

OpenStudy (lgbasallote):

7 students rated excellent there are a total of 7+14+ 4 = 25 students so the percent woulf be \[\frac{7}{25} \times 100\]

OpenStudy (anonymous):

this was a ACT practice question that I got incorrect. Can you please help me and explain to me step by step what to do on this problem?

OpenStudy (lgbasallote):

first identify the question...what is the question?

OpenStudy (anonymous):

what percent of the students thought the lecturer was excellent?

OpenStudy (lgbasallote):

so second step...how many students thought the lecturer was excellent?

OpenStudy (anonymous):

7

OpenStudy (lgbasallote):

good now percentage means \[\LARGE \frac{\text{number of students who thought teacher was excellent}}{\text{total number of students}} \times 100\] so what is the total number of students?

OpenStudy (anonymous):

25

OpenStudy (lgbasallote):

so substituting we have \[\huge \frac{7}{25} \times 100\] do you get that?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

that equals 28

OpenStudy (lgbasallote):

yup so 28% thought the teacher was excellent

OpenStudy (anonymous):

thank you so much!

OpenStudy (lgbasallote):

<tips hat>